# Towards Safer Pretraining: Analyzing and Filtering Harmful Content in Webscale datasets for Responsible LLMs Accepted at IJCAI 2025 Main Conference

### v2

@@ Abstract
Large language models (LLMs) have become integral to various real-world applications, leveraging massive, web-sourced datasets like Common Crawl, C4, and FineWeb for pretraining. While these datasets provide linguistic data essential for high-quality natural language generation, they often contain harmful content, such as hate speech, misinformation, and biased narratives. Training LLMs on such unfiltered data risks perpetuating toxic behaviors, spreading misinformation, and amplifying societal biases which can undermine trust in LLM-driven applications and raise ethical concerns about their use. This paper presents a large-scale analysis of inappropriate content across these datasets, offering a comprehensive taxonomy that categorizes harmful webpages into Topical and Toxic based on their intent. We also introduce a prompt evaluation dataset, a high-accuracy Topical and Toxic Prompt (TTP), and a transformer-based model (HarmFormer) for harmful content filtering. Additionally, we create a new multi-harm open-ended toxicity benchmark (HAVOC) and provide crucial insights into how models respond to adversarial toxic inputs. We share TTP, TTP-Eval, HAVOC and a sample of C4 inferenced on HarmFormer here  [^1] https://github.com/themendu/TowardsSaferPretraining. Our work offers insights into ensuring safer LLM pretraining and serves as a resource for Responsible AI (RAI) compliance.
@@ 1 Introduction
Large language models (LLMs) Achiam et al. (2023), Bubeck et al. (2023), Abdin et al. (2024), Dubey et al. (2024) derive their capabilities from pretraining on massive web-scale datasets such as Common CrawlRana (2010), C4 Raffel et al. (2020), and FineWeb Penedo et al. (2024). These datasets, while diverse, contain significant amounts of harmful content, hate speech, misinformation, and explicit material, which propagate into LLM output. For instance, as shown in Figure 1, prompting a Qwen2 model to generate self-harm methods yields dangerous, actionable suggestions. To mitigate such risks, practitioners rely on content moderation tools like keyword blocklists [^2] https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words, HateBERT Caselli et al. (2021), and Perspective API  [^3] https://perspectiveapi.com/. However, these tools suffer from critical limitations that undermine their effectiveness in real-world scenarios.
@@ 2 Related Works
Ideological Harm Objective analysis of conspiracy theories, investigative reporting on misinformation patterns, academic studies of societal biases. Example: www.bbcearth.com/news/why-some-people-believe-the-earth-is-flat Malicious dissemination of health disinformation, political radicalization content, pseudoscientific indoctrination. Example: www.reddit.com/…./genuine-questio-nwhy-are-conservatives-so-fucking
@@ 3.1 Taxonomy Structure
Our framework categorizes content through three complementary lenses: Safe (non-harmful material), Topical (contextual discussions requiring preservation), and Toxic (explicitly harmful content). As shown in Table 1, this tripartite structure is consistent across all categories of harm, allowing differential treatment of content types during the curation of the data sets. The Safe dimension serves as negative control space, containing material unrelated to predefined harms. Topical content encompasses contextually neutral discussions such as news reporting on violent events or academic analysis of conspiracy theories. Toxic material includes explicitly harmful content like violence promotion or self-harm instructions, aligned with established safety frameworks Inan et al. (2023).
@@ 4.1 TTP-Eval Set Creation
The prompt filtered the initial dataset down to approximately 50,000 documents for further analysis, serving as the candidate set for the next stages of labeling and GTTP-Eval set creation for evaluating the prompt. This approach allowed us to focus on the most relevant subset of documents, significantly reducing the dataset size while maintaining a high likelihood of capturing harmful content. The High Recall prompt is designed to identify the texts at a topic level, such as terrorism-related content within the Hate & Violence category, as well as its topical or toxic nature. To ensure high-quality annotations, we employed stratified sampling at the topic level in each of the harm to form a candidate set to create an eval-labeled dataset consisting of 491 web pages. We call this set TTP-Eval Set. These pages were carefully selected to represent all harm categories in our taxonomy, along with unrelated web pages to enhance diversity.
Each web page in our dataset was labeled by two in-house domain experts, both with over five years of experience in assessing content related to the identified harms, including Hate and Misinformation. The labeling process involved independent annotation by the two judges, followed by an agreement analysis. For web pages with disagreements, further discussion was conducted to refine the guidelines, ensuring consistency and clarity. The documents were manually labeled by domain experts, achieving a high inter-annotator agreement, as indicated by a Krippendorff’s Alpha score of 0.77. The distribution of various harms in our TTP-Eval set is presented in Table 2
@@ 4.2 TTP Prompt Creation
We developed our prompt (TTP) through an iterative process to improve the performance on the prompt eval dataset (TTP-Eval) on GPT-4 Omni. The prompt operates on a set of clear guidelines to ensure accurate classification, focusing on whether the thin line between simply discussing or actively promoting harmful content. Few-Shot prompting Brown et al. (2020), Chain of Thought Prompting Wei et al. (2022) helped us get better quality on the held out development set. The prompt’s performance on the test set is summarized in Table 4. The lower precision and recall of the prompt on Hate & Violence is explained by the ”Reporting webpages”. A reporting web page is a webpage purely reporting an incident of hate, or violence or any harm, with no intent of justifying or promoting the toxiciy included with the incident. These are those web pages that purely report hateful statements from a source against a receiver, with the source and the receiver being a person, organization or an entity.
@@
We benchmark Perspective API on our TTP-Evalset in table 4 to compare and highlight the quality and coverage of TTP prompt. Since the Perspective API is predominantly trained for hate, sexual and bias, it exhibits lower quality on our TTP-Evalset. To cope for the higher text lengths in TTP-Evalset, we chunk the texts into a length of 500 characters, and take the maximum score across chunks with 0.4 as the optimal threshold chosen from the dev set. We also run TTP on R1 DeepSeek-AI et al. (2025), Gemma 2 Team et al. (2024) and Gemini 2, and as expected the models perform poorly, as the prompt is tuned with GPT 4 Omni.
@@ 5.2 Model Training
Base Model Max Token Precision Recall F1
@@
Table 5: Toxic Content Detection Performance on our test set annotated by TTP. Precision, recall, and F1 scores compare base and fine-tuned (FT) models across aggregated toxic harm categories.
The detailed results of our experiments are presented in Table 5. As expected, Hate-BERT, which is primarily trained to detect hate speech, did not perform well on our test set (at a tuned threshold of 0.05). When finetuned to our scenario, Hate-BERT still underperforms with other models as it needs to relearn new harm categories along with the dimensions. We leverage Long Former’s (149M Parameters) 1024 context length capability over XLM-Roberta, which gave us an F1-Gain of 2 points over the same model trained with 512 context length. We finalize the architecture of the our HarmFormer model to be a LongFormer Beltagy et al. (2020), trained with 1024 context length across layers, a batch size of 16, learning rate of 2e-5, with AdamW Loshchilov and Hutter (2019) as the optimizer for 3 epochs over the train set.
@@ 6.1 Comparison with Open AI Moderation Dataset
To compare the robustness of TTP and lightweight model against SOTA models like Perspective API, and Llama Guard 3, we measure their performance on the publicly available Open AI Moderation Test set  [^4] https://github.com/openai/moderation-api-release as in Markov et al. (2023). This test set contains 1158 non toxic and 522 toxic text sentences varied across categories in Sexual, Hate, Violence and Self-Harm. As all the categories in the Open AI Moderation test set are inline and is a subset of toxic dimension of our taxonomy, we combined all of these together and reported the quality comparison in Table 7 for the single combined label.
We observe our prompt achieves a Precision and Recall of 0.72 and 0.91 respectively. We also test Llama Guard 3 Inan et al. (2023) in three different settings, the focused prompt setting which includes the categories the model was originally trained on, the Zero Shot setting, where we prompt the detailed description of the categories (provided by the dataset) to the language model, and Few-Shot, where we also append 2 examples per category along with the instructions. The HarmFormer and TTP still achieves a better performance compared to Perspective API and Llama-Guard 3, which proves their superiority to existing solutions in classifying text across multiple harms and text-lengths.
@@
The results in Table 7 show that TTP and HarmFormer are outperforming the Perspective and Llama Guard in toxic content classification on Open AI Moderation test set as well. This proves the effectiveness of our solutions across long and short form texts.
@@ 6.2.1 Toxic Content:
  • Common Crawl contains the highest proportion of toxic content (4.1%), driven by sexual content (2.2%), reflecting its unfiltered nature. Other toxic categories, including ideological harm (0.24%) and illegal activities (0.82%), are also prominent.
@@ 6.2.3 Category-Specific Trends in Toxic Content:
This analysis reveals that current LLM pretraining dataset curation practices fail to adequately distinguish harmful intent from socially critical or educational discourse. While filtering in C4 and FineWeb retains more topical content, keyword-based approaches often remove valuable context, resulting in gaps that risk propagating biases and toxicity in LLM training.
@@ 6.3 HAVOC: Harmful Abstractions and Violations in Open Completions Benchmark
To measure the toxicity generated by language models across input severity and harm categories, by creating a new benchmark of open ended prefixes. We sampled 10,376 snippets from web pages across the dimensions and harms identified by the TTP on the previously explained 3 Million candidate set in Section 5.1. A snippet consists of a group of sentences extracted from a web page using a prompt designed to capture content related to a known corresponding harm. We employed an additional prompt to determine the optimal break point between the snippet’s prefix and suffix, maximizing the risk of generating toxic content. We generate suffix from various LLMs given a prefix, aiming to study it’s toxic content generation in the text completion mode. Given that many of the web pages involve multiple harms, this multi-harm characteristic is also reflected in the snippets. Due to computational constraints, we restrict our benchmark size to 10,376. We call this newly created benchmark - Harmful Abstractions and Violations in Open Completions (HAVOC). In contrast to Real-Toxicity-Prompts (RTP)Gehman et al. (2020), HAVOC exhibits superiority with respect to the harm coverage (described in Figure 4) and the measurement quality (As shown in table 4).
We define a leak as a system-generated response that receives a Toxic label in any harm according to TTP. In the HAVOC benchmark, we apply the prompt to label both the input prefix and the combined prefix plus language model-generated suffix. A generation is classified as Provocative Leak if the prefix is labeled with Toxic for any harm and the combined prefix and generation also receive an Toxic label. Conversely, Passive Leak occurs when the prefix is labeled as either Safe or Topical, but the suffix is labeled as Toxic. Similarly, Neutral Leak is measured when the complete generation is labeled as Toxic, while the prefix remains labeled as Safe. In Table 9, we present some examples of our HAVOC benchmark. We hope HAVOC presents an alternative to quantify toxicity in language models.
@@
In 1987, Dr. Hans Hertel of Germany published his findings that microwaves cause changes to DNA (the genetic material) which can lead to cancer. He also found that microwaved foods have a lower nutritional value than non-microwaved foods. Topical This is because European women of all ages have a tendency to be even more reserved and do not like to show their feelings in public areas. Toxic
@@
We leverage Gemma 2 Team et al. (2024), Llama 3.2 Meta (2024), and Mistral V0.3 Mistral (2024) family of models to study their open ended generation on HAVOC and RTP. We use the base pretrained models for our evaluations. Due to hardware constraints, we leverage 2B, 9B & 27B model variants from Gemma 2, 1B & 3B variants from Llama 3.2, and use the 7B variant from the Mistral V0.3 family. We use Ollama Ollama (2024) to generate the completions for all the Language Models on our benchmark. We use greedy decoding to generate 200 output tokens from each prefix.
Figure 3: Model Averaged Neutral, Provocative and Passive Leakage comparison on HAVOC and RTP
@@
We aggregate the generations across the dataset and present our findings in Figure 3. As anticipated, Provocative Leak exhibits the highest percentage, driven by the presence of toxic content in the prefix prompt. The leakage percentages remain relatively consistent across different models, suggesting that there is potential for improving model training using non-toxic data. In table 10, we provide the model-averaged absolute leakage percentages for all six models. We report an overall leakage rate of 26.7% on the HAVOC benchmark. When challenged with a provocative input, a Language Model is expected to leak 76% of the time, raising key concerns. To have a holistic perspective at the harm level, we averaged the leakage percentage over all six models, and presented it over our 4 pivots in Figure 5. This highlights the problems over certain harms, with Sexual being the most problematic harm. Although significant, Hate and Ideological-Harm appear to be relatively less problematic compared to other harms. As shown in Figure 3, with the same experiment settings, HAVOC showcases double the leakage rates compared to RTP in Passive category, proving the effectiveness of the Topical Dimension in our taxonomy.

### v3

# Towards Safer Pretraining: Analyzing and Filtering Harmful Content in Webscale Datasets for Responsible LLMs This paper was accepted to the IJCAI 2025 Main Conference

@@ Abstract
Large language models (LLMs) have become integral to various real-world applications, leveraging massive, web-sourced datasets like Common Crawl, C4, and FineWeb for pretraining. While these datasets provide linguistic data essential for high-quality natural language generation, they often contain harmful content, such as hate speech, misinformation, and biased narratives. Training LLMs on such unfiltered data risks perpetuating toxic behaviors, spreading misinformation, and amplifying societal biases which can undermine trust in LLM-driven applications and raise ethical concerns about their use. This paper presents a large-scale analysis of inappropriate content across these datasets, offering a comprehensive taxonomy that categorizes harmful webpages into Topical and Toxic based on their intent. We also introduce a prompt evaluation dataset, a high-accuracy Topical and Toxic Prompt (TTP), and a transformer-based model (HarmFormer) for harmful content filtering. Additionally, we create a new multi-harm open-ended toxicity benchmark (HAVOC) and provide crucial insights into how models respond to adversarial toxic inputs. Our work offers insights into ensuring safer LLM pretraining and serves as a resource for Responsible AI (RAI) compliance.
@@ 1 Introduction
Large language models (LLMs)  Achiam et al. (2023); Bubeck et al. (2023); Abdin et al. (2024); Dubey et al. (2024) derive their capabilities from pretraining on massive web-scale datasets such as Common Crawl Rana (2010), C4 Raffel et al. (2020), and FineWeb  Penedo et al. (2024). These datasets, while diverse, contain significant amounts of harmful content, hate speech, misinformation, and explicit material, which propagate into LLM output. For instance, as shown in Figure 1, prompting a Qwen2 model to generate self-harm methods yields dangerous, actionable suggestions. To mitigate such risks, practitioners rely on content moderation tools like keyword blocklists [^1] https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words, HateBERT  Caselli et al. (2021), and Perspective API  [^2] https://perspectiveapi.com/. However, these tools suffer from critical limitations that undermine their effectiveness in real-world scenarios.  [^3] We share TTP, HAVOC and other sets here in Github.
@@ 2 Related Works
Ideological Harm Objective analysis of conspiracy theories, investigative reporting on misinformation patterns, academic studies of societal biases. Example: www.bbcearth.com/news/why-some-people-believe-the-earth-is-flat Malicious dissemination of health disinformation, political radicalization content, pseudoscientific indoctrination. Example: reddit.com/…./genuine-questio-nwhy-are-conservatives-so-fucking
@@ 3.1 Taxonomy Structure
Our framework categorizes content through three lenses: Safe (non-harmful material), Topical (contextual discussions requiring preservation), and Toxic (explicitly harmful content). As shown in Table 1, this tripartite structure is consistent across all categories of harm, allowing differential treatment of content types during the curation of the data sets. The Safe dimension serves as negative control space, containing material unrelated to predefined harms. Topical content encompasses contextually neutral discussions such as news reporting on violent events or academic analysis of conspiracy theories. Toxic material includes explicitly harmful content like violence promotion or self-harm instructions, aligned with established safety frameworks Inan et al. (2023).
@@ 4.1 TTP-Eval Set Creation
The prompt filtered the initial dataset down to approximately 50,000 documents for further analysis, serving as the candidate set for the next stages of labeling and TTP-Eval set creation for evaluating the prompt. This approach allowed us to focus on the most relevant subset of documents, significantly reducing the dataset size while maintaining a high likelihood of capturing harmful content. The High Recall prompt is designed to identify the texts at a topic level, such as terrorism-related content within the Hate & Violence category, as well as its topical or toxic nature. To ensure high-quality annotations, we employed stratified sampling at the topic level in each of the harm to form a candidate set to create an eval-labeled dataset consisting of 491 web pages. We call this set TTP-Eval Set. These pages were carefully selected to represent all harm categories in our taxonomy, along with unrelated web pages to enhance diversity.
Each web page in our dataset was labeled by two in-house domain experts, both with over five years of experience in judging the harms in our taxonomy including Hate and Misinformation. The labeling process involved independent annotation by the two judges, followed by an agreement analysis. For web pages with disagreements, further discussion was conducted to refine the guidelines, ensuring consistency and clarity. We observed a high inter-annotator agreement, as indicated by a Krippendorff’s Alpha score of 0.77. The distribution of various harms in our TTP-Eval set is presented in Table 2.
@@ 4.2 TTP Prompt Creation
We developed our prompt (TTP) through an iterative process to improve the performance on the prompt eval dataset (TTP-Eval). The prompt operates on a set of clear guidelines to ensure accurate classification, focusing on whether the thin line between simply discussing or actively promoting harmful content. Few-Shot prompting Brown et al. (2020), Chain of Thought Prompting Wei et al. (2022) helped us get better quality on the held out development set. The prompt’s performance on the test set is summarized in Table 4. The lower precision and recall of the prompt on Hate & Violence is explained by the ”Reporting webpages”, those that purely report an incident of hate, or violence or any harm, with no intent of justifying or promoting the toxicity. These are those web pages that purely report hateful statements from a source against a receiver, with the source and the receiver or the source being a person, organization or an entity.
@@
We benchmark Perspective API on our TTP-Evalset in table 4 to compare and highlight the quality and coverage of TTP prompt. Since the Perspective API is predominantly trained for hate, sexual and bias, it exhibits lower quality on our TTP-Evalset. To cope for the higher text lengths in TTP-Evalset, we chunk the texts into a length of 500 characters, and take the maximum score across chunks with 0.4 as the optimal threshold chosen from the dev set. We also run TTP on R1 DeepSeek-AI et al. (2025), Gemma 2 Gemma (2024) and Gemini 2, and as expected the models perform poorly, as the prompt is tuned with GPT 4 Omni.
@@ 5.2 Model Training
Base Model Tokens Precision Recall F1
@@
Table 5: Performance of base and fine-tuned (FT) models on our test set annotated by TTP on aggregated toxic harm categories.
Our HarmFormer model achieves an 85test set. The detailed results of our experiments are presented in Table 5. As expected, Hate-BERT, which is primarily trained to detect hate speech, did not perform well on our test set (at a tuned threshold of 0.05). When finetuned to our scenario, Hate-BERT still underperforms as it needs to relearn new harm categories along with the dimensions. We leverage Long Former’s 1024 context length capability over XLM-Roberta, which gave us an F1-Gain of 2 points over the same model trained with 512 context length. We finalize the architecture of the our HarmFormer model to be a LongFormer Beltagy et al. (2020), trained with 1024 context length across layers, a batch size of 16, learning rate of 2e-5, with AdamW Loshchilov and Hutter (2019) as the optimizer for 3 epochs over the train set.
@@ 6.1 Comparison with Open AI Moderation Dataset
To compare the robustness of TTP and lightweight model against Perspective API, and Llama Guard 3, we measure their performance on the publicly available Open AI Moderation Test set  [^4] https://github.com/openai/moderation-api-release as in Markov et al. (2023). This test set contains 1158 non toxic and 522 toxic text sentences varied across categories in Sexual, Hate, Violence and Self-Harm. Since Open AI Moderation set’s categories are a subset of out toxic dimension, we club these together into a single label and report the quality in Table 7.
We observe our prompt achieves a Precision and Recall of 0.72 and 0.91 respectively. We test Llama Guard 3 Inan et al. (2023) in three different settings, the focused prompt setting which includes the categories the model was originally trained on, the Zero Shot setting, where we prompt the detailed description of the categories (provided by the dataset) to the language model, and Few-Shot, where we also append 2 examples per category along with the instructions. The HarmFormer and TTP still achieves a better performance compared to Perspective API and Llama-Guard 3, which proves their superiority to existing solutions across text lengths.
@@
Table 7 shows that TTP and HarmFormer are outperforming the Perspective and Llama Guard on Open AI Moderation test set as well. This proves that they are better across long and short form texts.
@@ 6.2.1 Toxic Content:
  • Common Crawl contains the highest proportion of toxic content (4.1%), driven by sexual content (2.2%), reflecting its unfiltered nature. Other categories, including ideological harm (0.24%) and illegal activities (0.82%), are also prominent.
@@ 6.2.3 Category-Specific Trends in Toxic Content:
This analysis reveals that current pretraining dataset curation practices fail to adequately distinguish harmful intent from socially critical or educational discourse. While filtering in C4 and FineWeb retains more topical content, keyword-based approaches often remove valuable context resulting in gaps that risk propagating biases and toxicity in LLM training.
@@ 6.3 HAVOC: Harmful Abstractions and Violations in Open Completions Benchmark
To measure the toxicity generated by language models across input severity and harm categories, by creating a new benchmark of open ended prefixes. We sampled 10,376 snippets from web pages across the dimensions and harms identified by the TTP on the previously explained 3 Million candidate set in Section 5.1. A snippet consists of a group of sentences extracted from a web page using a prompt designed to capture content related to a known corresponding harm. We employed an additional prompt to determine the optimal break point between the snippet’s prefix and suffix, maximizing the risk of generating toxic content. Given that many of the web pages involve multiple harms, this multi-harm characteristic is also reflected in the snippets. Due to computational constraints, we restrict our benchmark size to 10,376. We call this newly created benchmark - Harmful Abstractions and Violations in Open Completions (HAVOC). In contrast to Real-Toxicity-Prompts (RTP)Gehman et al. (2020), HAVOC exhibits superiority with respect to the harm coverage (described in Figure 4) and the measurement quality (As shown in table 4).
We define a leak as a system-generated response that receives a Toxic label in any harm according to TTP. In the HAVOC benchmark, we apply the prompt to label both the input prefix and the combined prefix plus language model-generated suffix. A generation is classified as Provocative Leak if the prefix is labeled with Toxic for any harm and the combined prefix and generation also receive an Toxic label. Conversely, Passive Leak occurs when the prefix is labeled as Topical, but the suffix is labeled as Toxic. Similarly, Neutral Leak is measured when the complete generation is labeled as Toxic, while the prefix remains labeled as Safe. In Table 9, we present some examples of our HAVOC benchmark. We hope HAVOC presents an alternative to quantify toxicity in language models.
@@
He was thoughtful, responsive, open to questions, and in general didn’t have the Topical “I know everything” attitude that doctors can have. He also took time to explain things so I could understand them better. Toxic
@@
We leverage Gemma 2 Gemma (2024), Llama 3.2 Meta (2024), and Mistral V0.3 Mistral (2024) family of models to study their open ended generation on HAVOC and RTP. We use the base pretrained models for our evaluations. Due to hardware constraints, we leverage 2B, 9B & 27B model variants from Gemma 2, 1B & 3B variants from Llama 3.2, and use the 7B variant from the Mistral V0.3 family. We use Ollama Ollama (2024) to generate the completions for all the Language Models on our benchmark. We use greedy decoding to generate 200 output tokens from each prefix.
Figure 3: Model-averaged leakage (%) across Neutral, Passive, and Provocative tones on the HAVOC and RTP datasets.
@@
We aggregate the generations across the dataset and present our findings in Figure 3. As anticipated, Provocative Leak exhibits the highest percentage, driven by the presence of toxic content in the prefix prompt. The leakage percentages remain relatively consistent across different models, suggesting that there is potential for improving model training using non-toxic data. In table 10, we provide the model-averaged absolute leakage percentages for all six models. We report an overall leakage rate of 26.7% on the HAVOC benchmark. When challenged with a provocative input, an LLM is expected to leak 76% of the time. To have a holistic perspective at the harm level, we averaged the leakage percentage over all six models, and presented it in Figure 5. This figure highlights Sexual being the most problematic harm. Hate and Ideological-Harm appear to be relatively less problematic. As shown in Figure 3, with the same settings, HAVOC showcases double the leakage rates compared to RTP in Passive category, proving the effectiveness of the Topical Dimension in our taxonomy.